3. State how many protons , neutrons , and electrons there are in the atoms of the following isotopes: (a) 26He  (b)3 9Li  (c) 26 56Fe    (d) 92235 U


1
Expert's answer
2021-02-17T11:01:13-0500

(a) The atomic number (Z) of helium is 2 and the mass number (A) of the isotope is 6. Therefore, it has 2 protons, 2 electrons and N=A-Z=6-2=4 neutrons.

(b) The atomic number of lithium is 3 and the mass number of the isotope is 9. Therefore, it has 3 protons, 3 electrons and N=9-3=6 neutrons.

(c) The atomic number of iron is 26 and the mass number of the isotope is 56. Therefore, it has 26 protons, 26 electrons and N=56-26=30 neutrons.

(d) The atomic number of uranium is 92 and the mass number of the isotope is 235. Therefore, it has 92 protons, 92 electrons and N=235-92=143 neutrons.